Use point-slope form to determine the equation of the line passing through the two pointd given. Then write the equatioon in slope intercept form: (8,7)(1,-7)
a) y=2x-9
b) y=1/4x-9
c) y=1/2x-9
d) y=2x-7

Answer by nerdybill(7384)

Slope of:
(8,7)(1,-7)
m = (y2-y1)/(x2-x1)
m = (-7-7)/(1-8)
m = (-14)/(-7)
m = 2
.
Stuffing one of the points say (8,7) and the slope into the point-slope form:
y - y1 = m(x - x1)
y - 7 = 2(x - 8)
y - 7 = 2x - 16
y = 2x - 9